Use a calculator to solve the problem. Laura Shawhan has a pumpkin patch with 13 rows of pumpkin plants. There are 2 rows with 88 pumpkins each and 11 rows with 105 pumpkins each.How many pumpkins does Laura have?
Answer by checkley77(12844)

2*88+11*105=176+1155=1331 pumpkins total.